# ProMQA: Question Answering Dataset for Multimodal Procedural Activity Understanding

### v1

@@ 4.3 Results
In this experiment, we also obtained human performance as a comparison. We asked five first-stage annotators (§3.3) to solve 20 samples that they did not check in the verification process, given only recipes and video segments with questions. Table 8 shows that all the models we benchmark lag behind human performance, even the competitive proprietary models. Among the models, Claude 3.5 Sonnet performs relatively better than others, although the differences are somewhat marginal. In general, clean examples are easier for models than noisy examples, although the gaps vary depending on the models. Proprietary models are, on average, better on step-specific questions. Considering our results do not always align with those in the public benchmarks like MMMU and Video-MME, we believe our ProMQA can be complementary in evaluating models’ multimodal capabilities.
@@ 6 Related Work
QA as a task formulation is increasingly adopted for video QA datasets, e.g., NExT-QA (Xiao et al., 2021), EgoSchema (Mangalam et al., 2023), and Video-MME (Fu et al., 2024). While they are multimodal, i.e., a model takes video frames and a textual question as inputs, we argue that they are still rather video-oriented as only a short question consists of the textual part, compared to a pile of images from a video. For enhanced cross-modal comprehension, we present ProMQA where textual instructions are necessary to derive a correct answer in addition to a video and question (§2.1).
@@
Along with the advancement of LLMs, synthetic data generation is widely explored in various phases of model development, including pretraining (Gunasekar et al., 2023; Maini et al., 2024) and instruction tuning (Wang et al., 2023; Adler et al., 2024). Compared to those phases, it is underexplored in generating evaluation data with LLMs (Wu et al., 2024), possibly because of 1) the lack of quality assurance and 2) potential biases that it may introduce (Zheng et al., 2024; Panickssery et al., 2024). Addressing these challenges, we develop our ProMQA with additional human checks (§3.3), justified by the fairness-check experiments (§5).
@@ Appendix A ProMQA
Figure 5 illustrates the task formulation.777Icons in Figure 1, 3, 4, and 5 are from https://www.flaticon.com.
@@ A.1 External Knowldge
A.2 Human-written QAs
@@
A.3 Additional Statistics
In Figure 6 and 7, we show the counts of five starting words in questions and answers, sorted by question types.
@@ B.2 QA Generation
In the prompt exploration, the following are our findings: 1) Feeding the recipe as a whole hurts the approval rate compared to the target excerpt. This can be because the LLM needs to do extra reasoning to identify where to focus on a recipe. 2) Feeding the video as frames worsens the approval rate. Videos contain more information to contextualize the generation. However, the result suggests that even for the strong proprietary multimodal models, feeding information as text, if available, leads to better performance. In addition, feeding as frames costs is much less cost inefficient than feeing as text, as models require more tokens to process images.
@@ B.3 Verification
| Model | Avg. | w/ Error | Question Type | ||||||||
| clean | noisy | missing | next | order | measurement | preparation | technique | temperature | timing | ||
| Llama 3.1 8B Instruct | 25.7 | 25.9 | 25.6 | 35.1 | 22.6 | 25.0 | 5.9 | 0.0 | 20.0 | 16.0 | 14.3 |
| Llama 3.1 70B Instruct | 31.5 | 33.2 | 30.2 | 35.5 | 38.3 | 12.5 | 2.9 | 14.3 | 5.6 | 30.0 | 20.0 |
| LLaVA 1.5 7B (50f, 288p) & Llama 3.1 8B Instruct | 32.9 | 36.6 | 30.0 | 43.0 | 39.2 | 10.0 | 2.9 | 0.0 | 15.0 | 4.0 | 7.1 |
| LLaVA 1.5 13B (50f, 288p) & Llama 3.1 70B Instruct | 37.5 | 38.9 | 36.4 | 41.9 | 45.3 | 12.5 | 11.8 | 14.3 | 11.1 | 50.0 | 18.0 |
| VideoLLaMA2 7B (8f, 336p) | 39.3 | 45.7 | 34.2 | 47.5 | 47.3 | 22.5 | 0.0 | 0.0 | 40.0 | 8.0 | 14.3 |
| VideoLLaMA2 7B (16f, 336p) | 38.3 | 44.3 | 33.6 | 48.1 | 44.9 | 30.0 | 0.0 | 0.0 | 30.0 | 2.0 | 10.7 |
| VideoLLaMA2 72B (8f, 336p) | 39.8 | 49.4 | 32.2 | 46.3 | 49.7 | 20.0 | 0.0 | 21.4 | 11.1 | 25.0 | 8.0 |
| Qwen2 VL 7B (100f, 336p) | 33.8 | 38.6 | 30.0 | 43.4 | 36.8 | 30.0 | 0.0 | 0.0 | 30.0 | 6.0 | 14.3 |
| Qwen2-VL 72B (100f, 336p) | 31.2 | 32.1 | 30.4 | 34.1 | 37.0 | 45.0 | 0.0 | 10.7 | 0.0 | 20.0 | 14.0 |
| GPT-4o (50f, 765p) | 40.4 | 39.5 | 41.1 | 39.9 | 45.9 | 45.0 | 29.4 | 17.9 | 27.8 | 55.0 | 24.0 |
| GPT-4o (100f, 288p) | 38.9 | 39.2 | 38.7 | 37.2 | 44.0 | 37.5 | 32.4 | 21.4 | 22.2 | 50.0 | 34.0 |
| GPT-4o (250f, 288p) | 36.5 | 38.1 | 35.3 | 43.7 | 34.8 | 40.0 | 20.6 | 22.2 | 45.0 | 26.0 | 10.7 |
| Gemini (50f, 765p) | 25.2 | 27.0 | 23.8 | 27.4 | 29.7 | 15.0 | 17.6 | 7.1 | 16.7 | 20.0 | 12.0 |
| Gemini (100f, 288p) | 27.9 | 28.1 | 27.8 | 32.4 | 32.0 | 30.0 | 2.9 | 17.9 | 16.7 | 15.0 | 6.0 |
| Gemini (250f, 288p) | 27.7 | 30.4 | 25.6 | 30.1 | 31.8 | 32.5 | 8.8 | 22.2 | 10.0 | 8.0 | 25.0 |
| Claude (10f, 765p) | 44.1 | 48.9 | 40.4 | 44.6 | 58.2 | 27.5 | 8.8 | 14.3 | 5.6 | 25.0 | 28.0 |
| Claude (100f, 288p) | 36.8 | 43.8 | 31.3 | 48.4 | 37.5 | 22.5 | 14.7 | 16.7 | 25.0 | 12.0 | 10.7 |
| Human | 74.5 | 83.5 | 65.8 | — | — | — | — | — | — | — | — |
@@ C.2 Experimental Details
We use vllm for the inference of Llama 3.1 and LLaVA 1.5. For VideoLLaMA2888https://github.com/DAMO-NLP-SG/VideoLLaMA2 and Qwen2-VL.999https://github.com/QwenLM/Qwen2-VL, we follow their instructions to run their respective inference code. All the weights are downloaded from HuggingFace101010https://huggingface.co using transformers. We use 1~4 GPUs of A6000 (48GB), depending on the size of the models. Each inference took at most a few hours. For all proprietary models, we use their libraries to make API calls. Each prediction on all of our 401 examples in ProMQA costs 30~60 USD, depending on the model, the number of frames, and the resolution of each frame. All the results are based on a single run.
@@ C.3 Additional Results
Table 14 shows the additional benchmarking results by changing model size for open models and changing the number of frames and resolutions for proprietary models. The unimodal and Socratic models improve their performance as the sizes of their models increase. However, open multimodal models did not change the overall performance or even lowered their performance by a few points. As for proprietary models, under the fixed maximum input length, the number of frames trades off the resolutions. In our experiment, we found that higher resolution leads to better performance. However, the combinations we tried are rather limited, and there may exist a better combination, which we leave the exploration for future work. We also note here that we faced issues with limited maximum lengths with image-included prompts, compared to the ones listed on each API documentation or the ones when we tried text-only prompts. Presumably, this is due to the large file size of each image and the total data size of one input for each API request. We also leave it for future work on the workaround of how to feed many relatively high-resolution images in input prompts.

### v2

@@ 4.3 Results
In this experiment, we also obtained human performance as a comparison. We asked five first-stage annotators (§3.3) to solve 20 samples, out of 401 total examples, which they had not previously checked during the verification process. The sampling was done due to our budget. Performers were provided only recipes and video segments with questions. In Table 10, we provide the average performance, as well as the breakdown based on previous step types and question types. It shows that all the models we benchmark lag behind human performance, even the competitive proprietary models. Among the models, Claude 3.5 Sonnet performs relatively better than others, although the differences are somewhat marginal. In general, clean examples are easier for models than noisy examples, although the gap varies depending on each model. Proprietary models are, on average, better on step-specific questions. Additionally, we show the breakdown based on answer sources and types in Table 11. From the table, we can see that models generally perform better on examples with only machine-generated answers, although each model exhibits different preferences. Furthermore, we investigate the effect of answer counts of examples on performance. There is a weak common trend that models perform well on examples with a single answer and with 4 answers. Considering our results do not always align with those in the public benchmarks like Video-MME, we believe our ProMQA can be complementary in evaluating models’ multimodal capabilities.
@@ 6 Related Work
QA as a task formulation is increasingly adopted for video QA datasets, e.g., NExT-QA (Xiao et al., 2021), EgoSchema (Mangalam et al., 2023), OpenEQA (Majumdar et al., 2024), Video-MME (Fu et al., 2024), inter alia. While they are multimodal, i.e., a model takes video frames and a textual question as inputs, we argue that they are still rather video-oriented as only a short question consists of the textual part, compared to a pile of images from a video. While GazeVQA (Ilaslan et al., 2023) uniquely focuses on procedural tasks as QA, instructions are yet explicitly provided to systems, hence, only a short question with multiple choices and a video are the inputs. For enhanced cross-modal comprehension, we present ProMQA where textual instructions are necessary to derive a correct answer in addition to a video and question (§2.1).
@@
Along with the advancement of LLMs, synthetic data generation is widely explored in various phases of model development, including pretraining (Gunasekar et al., 2023; Maini et al., 2024) and instruction tuning (Wang et al., 2023; Adler et al., 2024). Compared to those phases, it is underexplored in generating evaluation data with LLMs (Wu et al., 2024), possibly because of the following two reasons: 1) The quality assurance is lacking, which can be mitigated by introducing multi-step machine and manual curation steps as EgoSchema. 2) Potential biases may be introduced (Zheng et al., 2024; Panickssery et al., 2024). Addressing these challenges, we develop our ProMQA with additional human checks (§3.3), justified by the fairness-check experiments (§5).
@@ 9 Ethical Consideration
Acknowledgements
We thank the anonymous reviewers for their insightful comments. This work is partially supported by Programs for Bridging the gap between R&D and the IDeal society (society 5.0) and Generating Economic and social value (BRIDGE) / Practical Global Research in the AI Robotics Services, implemented by the Cabinet Office, Government of Japan.
@@ Appendix A ProMQA
Figure 8 illustrates the task formulation.777Icons in Figure 1, 5, 6, and 8 are from https://www.flaticon.com.
@@ A.1 External Knowldge
A.2 Other Verification Criteria
Additionally, we ask annotators to check the naturalness and clarity of questions. Naturalness is to check if a question is natural/makes sense to ask. For instance, when a question like “Did I forget to do something before <stepX>?” is asked, people usually assume that <stepX> has already been passed (with or without errors). So, if the question is asked when <stepX> is yet to be performed, this question will be unnatural/nonsensical. This criterion filters out this type of nonsensical question. Clarity filters out vague/too general questions, especially questions asking about non-procedural aspects. For instance, a question like “What did I do wrong?” can target non-procedural errors, e.g., “Too many dishes are left in the sink.” or “The countertop is too messy.” which we encountered in our preliminary QA generation and benchmarking experiments. To focus on the procedural questions, we added this criterion.
A.3 Human-written QAs
@@
A.4 Additional Statistics
In Figure 9 and 10, we show the counts of five starting words in questions and answers, sorted by question types.
@@ A.4 Additional Statistics
A.5 Statistical Power Analysis
Following Card et al. (2020), we conduct their statistical power analysis to estimate the performance difference required to detect statistical significance between systems with confidence. We first estimate the baseline accuracy based on the performance of GPT-4o, and the agreement rate based on GPT-40 and Claude 3.5 Sonnet, . Given these numbers, the simulation-based analysis888https://github.com/dallascard/NLP-power-analysis shows that at least accuracy point difference would be needed to detect significance with confidence.
@@ B.2 QA Generation
In the prompt exploration, the following are our findings: 1) Feeding the recipe as a whole hurts the approval rate compared to the target excerpt. This can be because the LLM needs to do extra reasoning to identify where to focus on a recipe. 2) Feeding the video as frames worsens the approval rate. Videos contain more information to contextualize the generation. However, the result suggests that even for the strong proprietary multimodal models, feeding information as text, if available, leads to better performance. In addition, feeding as frames costs is much more expensive than feeding as text, as models require more tokens to process images.
@@ B.3 Verification
| Model | Avg. | w/ Error | Question Type | ||||||||||
| clean | noisy | missing | next | order | measurement | preparation | technique | temperature | timing | ||||
| Llama 3.1 8B Instruct | 25.7 | 25.9 | 25.6 | 35.1 | 22.6 | 25.0 | 5.9 | 0.0 | 20.0 | 16.0 | 14.3 | ||
| Llama 3.1 70B Instruct | 31.5 | 33.2 | 30.2 | 35.5 | 38.3 | 12.5 | 2.9 | 14.3 | 5.6 | 30.0 | 20.0 | ||
32.9 | 36.6 | 30.0 | 43.0 | 39.2 | 10.0 | 2.9 | 0.0 | 15.0 | 4.0 | 7.1 | ||
37.5 | 38.9 | 36.4 | 41.9 | 45.3 | 12.5 | 11.8 | 14.3 | 11.1 | 50.0 | 18.0 | ||
| VideoLLaMA2 7B (8f, 336p) | 39.3 | 45.7 | 34.2 | 47.5 | 47.3 | 22.5 | 0.0 | 0.0 | 40.0 | 8.0 | 14.3 | ||
| VideoLLaMA2 7B (16f, 336p) | 38.3 | 44.3 | 33.6 | 48.1 | 44.9 | 30.0 | 0.0 | 0.0 | 30.0 | 2.0 | 10.7 | ||
| VideoLLaMA2 72B (8f, 336p) | 39.8 | 49.4 | 32.2 | 46.3 | 49.7 | 20.0 | 0.0 | 21.4 | 11.1 | 25.0 | 8.0 | ||
| Qwen2 VL 7B (100f, 336p) | 33.8 | 38.6 | 30.0 | 43.4 | 36.8 | 30.0 | 0.0 | 0.0 | 30.0 | 6.0 | 14.3 | ||
| Qwen2-VL 72B (100f, 336p) | 31.2 | 32.1 | 30.4 | 34.1 | 37.0 | 45.0 | 0.0 | 10.7 | 0.0 | 20.0 | 14.0 | ||
| GPT-4o (50f, 765p) | 40.4 | 39.5 | 41.1 | 39.9 | 45.9 | 45.0 | 29.4 | 17.9 | 27.8 | 55.0 | 24.0 | ||
| GPT-4o (100f, 288p) | 38.9 | 39.2 | 38.7 | 37.2 | 44.0 | 37.5 | 32.4 | 21.4 | 22.2 | 50.0 | 34.0 | ||
| GPT-4o (250f, 288p) | 36.5 | 38.1 | 35.3 | 43.7 | 34.8 | 40.0 | 20.6 | 22.2 | 45.0 | 26.0 | 10.7 | ||
| Gemini 1.5 Pro (50f, 765p) | 25.2 | 27.0 | 23.8 | 27.4 | 29.7 | 15.0 | 17.6 | 7.1 | 16.7 | 20.0 | 12.0 | ||
| Gemini 1.5 Pro (100f, 288p) | 27.9 | 28.1 | 27.8 | 32.4 | 32.0 | 30.0 | 2.9 | 17.9 | 16.7 | 15.0 | 6.0 | ||
| Gemini 1.5 Pro (250f, 288p) | 27.7 | 30.4 | 25.6 | 30.1 | 31.8 | 32.5 | 8.8 | 22.2 | 10.0 | 8.0 | 25.0 | ||
| Claude 3.5 Sonnet (10f, 765p) | 44.1 | 48.9 | 40.4 | 44.6 | 58.2 | 27.5 | 8.8 | 14.3 | 5.6 | 25.0 | 28.0 | ||
| Claude 3.5 Sonnet (100f, 288p) | 36.8 | 43.8 | 31.3 | 48.4 | 37.5 | 22.5 | 14.7 | 16.7 | 25.0 | 12.0 | 10.7 | ||
| Human | 74.5 | 83.5 | 65.8 | — | — | — | — | — | — | — | — | ||
| Model | Avg. | Answer Source | Answer Type | ||||||||
| machine | human | both | direct | direct & suggestion | direct & intervention | suggestion | all | ||||
| Llama 3.1 8B Instruct | 25.7 | 23.4 | 22.0 | 33.3 | 26.4 | 30.8 | 9.1 | 6.2 | 100.0 | ||
| Llama 3.1 70B Instruct | 31.5 | 33.4 | 28.0 | 30.5 | 31.7 | 30.8 | 34.1 | 25.0 | 0.0 | ||
32.9 | 34.1 | 29.9 | 32.9 | 33.2 | 37.2 | 27.3 | 6.2 | 100.0 | ||
37.5 | 40.7 | 32.9 | 34.8 | 38.8 | 34.6 | 34.1 | 12.5 | 0.0 | ||
| VideoLLaMA2 7B (8f, 336p) | 39.3 | 43.7 | 31.7 | 36.2 | 40.9 | 39.7 | 25.0 | 12.5 | 0.0 | ||
| VideoLLaMA2 7B (16f, 336p) | 38.3 | 41.8 | 32.3 | 35.7 | 39.4 | 41.0 | 27.3 | 12.5 | 0.0 | ||
| VideoLLaMA2 72B (8f, 336p) | 39.8 | 45.1 | 32.9 | 34.3 | 41.7 | 34.6 | 31.8 | 12.5 | 0.0 | ||
| Qwen2 VL 7B (100f, 336p) | 33.8 | 36.7 | 28.0 | 32.4 | 35.3 | 28.2 | 29.5 | 12.5 | 0.0 | ||
| Qwen2-VL 72B (100f, 336p) | 31.2 | 35.0 | 22.0 | 30.5 | 30.8 | 33.3 | 27.3 | 37.5 | 100.0 | ||
| GPT-4o (50f, 765p) | 40.4 | 41.4 | 29.3 | 47.1 | 40.2 | 39.7 | 52.3 | 18.7 | 50.0 | ||
| GPT-4o (100f, 288p) | 38.9 | 38.8 | 27.4 | 48.1 | 38.1 | 38.5 | 56.8 | 31.2 | 0.0 | ||
| GPT-4o (250f, 288p) | 36.5 | 36.7 | 26.8 | 43.8 | 36.0 | 39.7 | 45.5 | 25.0 | 0.0 | ||
| Gemini 1.5 Pro (50f, 765p) | 25.2 | 27.1 | 18.3 | 26.7 | 24.8 | 25.6 | 34.1 | 12.5 | 50.0 | ||
| Gemini 1.5 Pro (100f, 288p) | 27.9 | 28.3 | 20.7 | 32.9 | 27.6 | 34.6 | 22.7 | 25.0 | 0.0 | ||
| Gemini 1.5 Pro (250f, 288p) | 27.7 | 26.6 | 28.0 | 29.5 | 27.9 | 29.5 | 20.5 | 31.2 | 0.0 | ||
| Claude 3.5 Sonnet (10f, 765p) | 44.1 | 48.1 | 35.4 | 42.9 | 45.0 | 42.3 | 43.2 | 25.0 | 0.0 | ||
| Claude 3.5 Sonnet (100f, 288p) | 36.8 | 38.8 | 30.5 | 37.6 | 36.6 | 42.3 | 38.6 | 18.7 | 0.0 | ||
@@ C.2 Experimental Details
We use vllm for the inference of Llama 3.1 and LLaVA 1.5. For VideoLLaMA2999https://github.com/DAMO-NLP-SG/VideoLLaMA2 and Qwen2-VL.101010https://github.com/QwenLM/Qwen2-VL, we follow their instructions to run their respective inference code. All the weights are downloaded from HuggingFace111111https://huggingface.co using transformers. We use 1~4 GPUs of A6000 (48GB), depending on the size of the models. Each inference took at most a few hours. For all proprietary models, we use their libraries to make API calls. Each prediction on all of our 401 examples in ProMQA costs 30~60 USD, depending on the model, the number of frames, and the resolution of each frame. All the results are based on a single run.
@@ C.3 Additional Results
Table 15 and 16 show the additional benchmarking results by changing model size for open models and changing the number of frames and resolutions for proprietary models. The unimodal and Socratic models improve their performance as the sizes of their models increase. However, open multimodal models did not change the overall performance or even lowered their performance by a few points. As for proprietary models, under the fixed maximum input length, the number of frames trades off the resolutions. In our experiment, we found that higher resolution leads to better performance. However, the combinations we tried are rather limited, and there may exist a better combination, which we leave the exploration for future work. We also note here that we faced issues with limited maximum lengths with image-included prompts, compared to the ones listed on each API documentation or the ones when we tried text-only prompts. Presumably, this is due to the large file size of each image and the total data size of one input for each API request. We also leave it for future work on the workaround of how to feed many relatively high-resolution images in input prompts.